x+2y=23
5x+10y=55

Answer by mananth(16949)

1 x + 2 y = 23 .............1
5 x + 10 y = 55 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 1
-5 x -10 y = -115
5 x 10 y = 55
Add the two equations
0 x = -60
this is absurd
so there is no solution to this system. If you draw a graph you will seee that they are parallel lines